### v1

# RoMedQA: The First Benchmark for Romanian Medical Question Answering

Abstract
Question answering (QA) is an actively studied topic, being a core natural language processing (NLP) task that needs to be addressed before achieving Artificial General Intelligence (AGI). However, the lack of QA datasets in specific domains and languages hinders the development of robust AI models able to generalize across various domains and languages. To this end, we introduce RoMedQA, the first Romanian QA benchmark for the medical domain, alongside a comprehensive evaluation of state-of-the-art large language models (LLMs). We construct a high-quality and large-scale dataset comprising 102,646 QA pairs related to cancer patients. The questions regard medical case summaries of 1,011 patients, requiring either keyword extraction or reasoning to be answered correctly. RoMedQA is the result of a time-consuming manual annotation process carried out by seven physicians specialized in oncology or radiotherapy, who spent a total of about 2,100 work hours to generate the QA pairs. We experiment with four LLMs from distinct families of models on RoMedQA. Each model is employed in two scenarios, namely one based on zero-shot prompting and one based on supervised fine-tuning. Our results show that fine-tuned models significantly outperform their zero-shot counterparts, clearly indicating that pretrained models fail to generalize on RoMedQA. Our findings demonstrate the importance of both domain-specific and language-specific fine-tuning for reliable clinical QA in Romanian. We publicly release our dataset and code at https://github.com/ana-rogoz/RoMedQA.
Figure 1: An image illustrating our dataset creation and model benchmarking stages.
1 Introduction
The development of robust question answering (QA) systems is an important goal in natural language processing (NLP), representing one of prerequisites for building systems that reach human-level intelligence. Despite the significant advances in the area of Large Language Models (LLMs) Pal and Sankarasubbu (2024); Touvron et al. (2023); Jiang et al. (2023); Taori et al. (2023); Köpf et al. (2023), which are able to perform remarkably well across various QA tasks Brown et al. (2020); Chowdhery et al. (2023); Rajpurkar et al. (2016); Hendrycks et al. (2021); Achiam et al. (2023); Singhal et al. (2023, 2025), QA remains a challenging task within specialized domains, such as the medical field, or in low-resource languages, such as Romanian. For instance, the task of medical QA comes with unique challenges due to the complexity of clinical narratives and the necessity for deep understanding of medical terms. Moreover, there is a soaring demand for high accuracy, as mistakes can have serious consequences in daily medical practice assisted by AI agents. While large-scale datasets such as PubMedQA Jin et al. (2019), emrQA Pampari et al. (2018), and MedQA Jin et al. (2021) have driven significant progress in English medical QA, the availability of similar resources in other languages, especially underrepresented ones, remains limited. To the best of our knowledge, for the Romanian language, there are no prior efforts in the realm of QA resources for the medical domain.
To address this gap, we introduce RoMedQA, the first benchmark for Romanian Medical Question Answering. As illustrated in Figure 1, the benchmark creation involves two phases, (i) manual data collection and annotation, and (ii) model benchmarking. The dataset comprises 102,646 high-quality QA pairs from real-world clinical records of 1,011 oncology patients (796 patients with breast cancer and 215 patients with lung cancer). The QA pairs are the results of a manual annotation process carried out by physicians specialized in oncology and radiotherapy, who collectively spent nearly 2,100 hours to generate the pairs, ensuring both linguistic quality and medical correctness. RoMedQA includes 76,416 QA pairs about breast cancer patients and 26,230 about lung cancer patients, with questions grounded in medical case summaries (epicrises).
Dataset Language Domain # QA pairs
SQuAD Rajpurkar et al. (2016) English General knowledge 107,785
Natural Questions Kwiatkowski et al. (2019) English General knowledge 307,373
XQuAD Artetxe et al. (2020) Multilingual (11 languages) General knowledge 1,190
MLQA Lewis et al. (2020) Multilingual (7 languages) General knowledge 17,000
TyDi QA Clark et al. (2020) Multilingual (11 languages) Information-seeking 204,000
PubMedQA Jin et al. (2019) English Medical 212,300
emrQA Pampari et al. (2018) English Medical 455,837
MedQA Jin et al. (2021) English Medical exams 12,723
MedMCQA Pal et al. (2022) English Medical exams (India) 193,100
JuRo Crăciun et al. (2025) Romanian Legal 10,836
RoITD Nicolae and Tufiş (2021) Romanian Technical 9,500
LiRo / XQuAD-Ro Dumitrescu et al. (2021) Romanian General knowledge 1,190
RoMedQA (ours) Romanian Medical 102,646
Table 1: Comparison between RoMedQA and existing QA datasets across three dimensions, namely language coverage, domain focus, and dataset size. Existing QA datasets range from general-purpose English resources to multilingual datasets and domain-specific collections in areas such as legal, technical and medical fields. RoMedQA is the first benchmark for medical QA in Romanian.
To prevent data leakage across splits and support fair model development and evaluation, we provide official training, validation, and test splits at the patient level, i.e. a medical record can appear in only one of the splits. We benchmark four models on RoMedQA, namely two LLMs specialized on Romanian (RoLLaMA2-7B and RoMistral-7B) Masala et al. (2024), a long-context LLM (Phi-4-mini-instruct) Abouelenin et al. (2025), and an LLM tuned on biomedical data (LLaMA3-OpenBioLLM-8B) Pal and Sankarasubbu (2024). Each model is assessed in both zero-shot and supervised fine-tuning setups, while testing various prompt formats. Our results show that all LLMs significantly benefit from fine-tuning, whereas zero-shot variants perform poorly. These findings emphasize the critical importance of both domain and language adaptation for reliable clinical QA in Romanian. Among all evaluated models, the fine-tuned Phi-4-mini-instruct achieves the best performance, significantly outperforming other approaches. However, the best-performing model achieves an F1 score of only $0.667$, suggesting that RoMedQA is a challenging dataset, highlighting the need for the development of more robust models, capable of better generalizing to specialized domains and low-resource languages.
In summary, our contributions are fourfold:
  • We introduce RoMedQA, the first and only large-scale medical QA dataset in Romanian, comprising 102,646 QA instances obtained with the help of trained physicians.
  • We benchmark four LLMs from distinct families of models across multiple configurations and report comprehensive performance metrics under both zero-shot and fine-tuning configurations.
  • We demonstrate that domain-specific and language-specific adaptation is crucial for reliable clinical QA in low-resource languages such as Romanian.
  • We release the dataset and code to reproduce the results in a public repository, where patient sensitive data is fully anonymized.
2 Related Work
English and multi-lingual QA datasets. Question answering emerged as a fundamental task in NLP, with several datasets being developed to support research across various domains and languages. Popular QA datasets focusing on the English language, such as SQuAD Rajpurkar et al. (2016) and Natural Questions Kwiatkowski et al. (2019), established the foundation for extractive and generative QA systems. These datasets showed the importance of large-scale high-quality annotations for training robust QA models. Later on, given the need for cross-language understanding, more and more datasets covering several languages started to arise. Some notable multilingual efforts include XQuAD Artetxe et al. (2020), which extends SQuAD to 11 languages, and MLQA Lewis et al. (2020), which provides QA datasets in seven languages with cross-lingual evaluation capabilities. Furthermore, TyDi QA Clark et al. (2020) advanced multilingual QA by covering structurally different languages and focusing on information-seeking questions. However, these efforts did not include Romanian among the target languages.
Medical QA datasets. Medical QA represents a challenging domain that requires deep understanding of a broad range of aspects, such as clinical terminology and medical reasoning. Several English medical QA datasets have been already developed to tackle the particularities of the medical domain. For instance, MedQA Jin et al. (2021) provides over 12,000 multiple-choice questions from medical licensing exams across three countries, testing complex clinical reasoning abilities. Additionally, the emrQA dataset Pampari et al. (2018) introduces one of the largest QA medical resources with over 455,000 question-answer pairs automatically generated from electronic medical records. PubMedQA Jin et al. (2019) focuses on biomedical literature comprehension, with questions derived from research article titles and abstracts. MedMCQA Pal et al. (2022) expands medical QA to include Indian medical entrance exams, with 193,100 questions across 21 medical subjects, demonstrating the importance of diverse medical knowledge sources. Despite its data source, the QA pairs contained by MedMCQA are in English.
In contrast to the significant advances in English medical QA, the landscape of non-English medical QA datasets remains scarce. To the best of our knowledge, we are the first to introduce a Romanian QA dataset specifically tailored for the medical domain. RoMedQA contains original clinically-grounded content rather than translations, addressing a critical gap in current literature.
Romanian QA datasets. For the Romanian language, a few QA datasets have been independently developed. Two such datasets focus on domain specific understanding. One is the JuRo dataset Crăciun et al. (2025), which provides a collection of Romanian juridical texts with associated questions, and the other is the RoITD dataset Nicolae and Tufiş (2021), which addresses technical domain questions in the IT sector. Additionally, the LiRo benchmark Dumitrescu et al. (2021), which provides a comprehensive evaluation platform for Romanian NLP tasks, contributed with a subset of XQuAD Artetxe et al. (2020) for Romanian, consisting of translated versions of English questions. None of the existing datasets for Romanian QA is tailored for the medical domain. Moreover, existing datasets are typically small (containing between 1,000 and 11,000 QA pairs). With 102,646 QA pairs, RoMedQA is one order of magnitude larger than the largest dataset for Romanian QA.
RoMedQA vs. QA dataset landscape. In Table 1, we present a comprehensive comparison between RoMedQA and existing QA datasets across different languages and domains. RoMedQA fills in a gap in the landscape of medical QA resources by providing the first large-scale Romanian medical QA dataset with 102,646 question-answer pairs focused on oncology. While other Romanian QA datasets exist for other domains, none of them specifically addresses the medical domain with clinically-relevant content in Romanian. This positions RoMedQA as the first and only resource for advancing medical NLP capabilities for Romanian, an underrepresented language in the QA field.
Number of patients Number of samples
Train Validation Test Total Train Validation Test Total
Breast cancer 557 119 120 796 53,472 11,424 11,520 76,416
Lung cancer 150 32 33 215 18,300 3,904 4,026 26230
Total 707 151 153 1,011 71,772 15,328 15,546 102,646
Table 2: A summary of the number of patients and corresponding QA pairs included in the RoMedQA benchmark across two types of cancer, breast and lung. The splits are performed at the patient level, into training (70%), validation (15%), and test (15%), to prevent information leakage across splits. We report the number of patients per cancer type per split, and the corresponding number of QA samples generated for each split. The final row aggregates the total number of unique patients and QA pairs across both cancer types for each split. RoMedQA comprises a total 102,646 QA pairs for 1,011 patients.
3 Dataset
3.1 Data Collection
RoMedQA is obtained through the manual extraction of information from medical documents, namely medical summaries of patients registered in the Oncology Department of Colţea Clinical Hospital in Bucharest, Romania. The epicrises, originally drafted in Word format by the physicians attending each patient, were manually reviewed by a group of seven physicians involved in constructing RoMedQA. In this group, two experienced physicians, who led the data collection process, established the most prevalent cancer sites in their hospital, namely breast and lung. They selected patients with either one of these diagnostics, who received treatment at the hospital between 2019 and 2024. This selection process led to a set of 796 patients with breast cancer and 215 patients with lung cancer.
The two experienced physicians further proposed a set of reference questions for each cancer site, comprising 48 questions for breast cancer and 61 questions for lung cancer. The other five physicians proposed 8 to 10 equivalent reformulations of the reference questions. Next, the physicians proceeded by extracting the answers for each unique question from each epicrisis and entering the respective answers into a database.
Data extraction and completion were carried out via a two-level process. The first level was performed by five medical oncology and radiotherapy residents, who meticulously completed the database of QA pairs following prior training based on the annotation guidelines provided by the experienced physicians. The second level was carried out by two board-certified medical oncology specialists, holders of a good standing certificate issued by the Romanian College of Physicians and with proven experience in good medical and clinical research practices. They verified the accuracy and consistency of the data entered by the residents. In total, the annotation process involved seven physicians, with a cumulative effort of approximately 2,100 hours (equivalent to about 300 hours per physician). The number of patients processed by each resident ranged from a minimum of 200 to a maximum of 220 cases, depending on individual availability and case complexity. To measure the inter-annotator agreement, a number of 350 QA pairs were annotated by all five physicians. The average Cohen’s $\kappa$ score between each pair of annotators is $0.9562$, which indicates an almost perfect agreement.
Furthermore, each experienced specialist meticulously reviewed the QA pairs associated with about 500-510 patients. This distribution of the workload was planned to ensure a relatively balanced annotation process. For data extraction, the experienced physicians developed an annotation guideline, establishing clear completion rules, including: staging of cancers according to the Tumor Node Metastasis (TNM) classification; grading of adverse events according to the Common Terminology Criteria for Adverse Events (CTCAE) system; standardization of medical abbreviations (e.g. “CHT” for chemotherapy, “RT” for radiotherapy, “NK” for “not known” values); handling incomplete information. In addition to these general rules, the guideline included a detailed list of answer options to be extracted and recorded in a standardized format: Eastern Cooperative Oncology Group (ECOG) performance status, menopausal status, molecular subtype, tumor proliferation index (Ki-67), histological grade, tumor size, family history of cancer, personal history of other cancers, body mass index (BMI), tumor markers, presence/absence of disease progression, number of metastatic sites, metastatic locations, death, etc.
Regarding research ethics, we emphasize that all data was fully anonymized prior to processing by removing any direct identifiers (names, dates of birth, personal identification numbers, etc.). The use of anonymized data for research purposes is based on the informed consent of patients, as recorded in the medical chart of Colţea Clinical Hospital and signed individually by each patient at admission. The database was approved by the Ethics Committee of Colţea Clinical Hospital where the research was conducted, in accordance with Decision No. 19091, dated October 5th, 2021. The entire process complies with the requirements of the General Data Protection Regulation (GDPR), the principles of the Declaration of Helsinki, and the standards of good medical and research practices.
3.2 Preprocessing
To support model training and evaluation, we constructed two domain-specific datasets targeting question answering in oncology given a clinical context: one for breast cancer and one for lung cancer.
Breast cancer. We collected data from 796 patients, each associated with a complete epicrisis, i.e. a detailed clinical summary for each patient, and a corresponding set of answers to 48 unique medical questions. To simulate natural variation in clinical queries, each question is reformulated by multiple physicians, leading multiple alternative formulations. For every patient-question pair, we randomly sample two question alternatives and generate two QA instances, each structured as a tuple of the form (patient ID, epicrisis, question, answer). This process resulted in a total of 76,416 unique examples.
Lung cancer. We compiled data from another 215 patients, each accompanied by an epicrisis summarizing their clinical history. For each patient, we curated a set of answers to 61 unique and clinically relevant questions. Similar to the breast cancer setup, each question is manually paraphrased into multiple linguistically diverse formulations to better reflect natural question variability. For each patient-question pair, we randomly sampled two different question variants. This leads to a total of 26,230 unique QA instances, capturing both clinical diversity and linguistic variation.
3.3 Data Partitioning and Statistics
We split the dataset into training, validation and test, using a $70\%/15\%/15\%$ split ratio, ensuring that the division occurs at the patient level. We prefer this split over splitting individual QA pairs because it guarantees that all the data from a single patient, including their epicrisis and associated QA pairs, remains within exactly one subset, thereby preventing data leakage among splits.
We present the distribution of our dataset across training, validation, and test sets in Table 2. For the breast cancer subset, our split results in 557 patients allocated for training, 119 for validation, and 120 for testing. In terms of QA pairs, the split distributes 53,472 samples for training, 11,424 for validation, and 11,520 for testing, respectively. Similarly, the lung cancer subset is divided into 150 training, 32 validation, and 33 test patients, yielding 18,300, 3,904, and 4,026 QA pairs, respectively. The splits are independently generated for each cancer type. We next combine the corresponding partitions to create our final unified dataset, called RoMedQA. In total, RoMedQA comprises 1,011 unique patients (796 suffering from breast cancer, and 215 from lung cancer) and 102,646 question-answer pairs, providing a substantial resource for training and evaluating Romanian medical QA systems.
In Table 3, we report statistics about the number of tokens in epicrises, questions and answers included RoMedQA. Epicrises range from 207 to 34,247 tokens, with a mean value of 7,171 tokens, where shorter documents typically correspond to patients under investigation or at early treatment stages, while longer ones capture extensive treatment histories with multiple interventions. This variation can pose challenges for some models with limited context windows. While epicrises are typically lengthy, questions and answers are much shorter. Questions maintain relative consistency, between 7 and 30 tokens, with a mean value of 15 tokens, reflecting the consistency of clinical query formulations. The answers are represented by predominantly brief texts with a mean of 4 tokens, varying in length from one to 442 tokens. Longer responses are typical for questions involving complex medical reasoning.
Data type Number of tokens
Min Max Mean Total
Epicrisis 207 34,247 7,171 7,249,881
Question 7 30 15 1,621,561
Answer 1 442 4 420,708
Table 3: A summary of the minimum, maximum, average and total number of tokens for the three kinds of data (epicrisis, question and answer) available in RoMedQA.
Figure 2: Distribution of question types in the RoMedQA dataset. The dataset comprises three main categories: binary (yes/no) questions, extractive questions (answers explicitly found in the epicrisis), and reasoning questions (requiring inference beyond explicit mentions). Percentages are computed over the full set of 102,646 QA pairs. Best viewed in color.
We categorize the questions into three types: (i) binary (yes/no); (ii) extractive, where the answer appears explicitly in the clinical context (epicrisis); and (iii) reasoning, where the answer must be inferred from multiple clues within the epicrisis. The distribution of these categories in RoMedQA is shown in Figure 2. Within the binary subset, the label distribution is moderately skewed, with roughly two thirds of answers being “Nu” (“No”) and one third being “Da” (“Yes”). The slightly skewed distribution resulted naturally from the data collection process.
4 Methods
To address the task of medical question answering in Romanian, we consider a diverse set of LLMs and use them under various configurations, considering both zero-shot and fine-tuning settings. Our goal is to understand how distinct architectural specializations (e.g. language-specific, domain-specific, or long-context models) and alternative prompt formulations affect performance on the Romanian clinical QA task. We select four distinct LLM families. Two of them are pretrained on Romanian, namely RoLLaMA2-7B Masala et al. (2024), which is adapted from LLaMA 2 Touvron et al. (2023), and RoMistral-7B Masala et al. (2024), which is adapted from the Mistral architecture Jiang et al. (2023). The third model, namely Phi-4-mini-instruct Abouelenin et al. (2025), is optimized for long-context inputs. This is useful for the processing of entire epicrises, which can often get to be very long (see Table 3), especially for patients with a long medical history. Finally, given the target domain, the fourth model, LLaMA3-OpenBioLLM-8B Pal and Sankarasubbu (2024), is pretrained on biomedical corpora.
All four LLMs are employed in two setups, zero-shot prompting and end-to-end fine-tuning. Due to the large number of parameters in the chosen models, we adopt a parameter-efficient fine-tuning strategy based on Low-Rank Adaptation (LoRA) Hu et al. (2022) specialized on the causal language modeling task. By comparing the results obtained in the two setups, we can precisely contextualize the impact of fine-tuning on RoMedQA. In addition, we also test two alternative prompts, where the order between context (epicrisis) and question is interchanged. Finally, to determine the impact of zero-shot prompting, we include a random token selector and a majority answer model as baselines for the QA task. Next, we describe all evaluated methods in detail.
4.1 Romanian-Adapted LLMs
We start from two instruction-tuned language models specifically adapted for Romanian, RoLLaMA2-7B-Instruct and RoMistral-7B-Instruct. Both models are part of the recent efforts to build high-quality Romanian LLMs by further pretraining on Romanian corpora and applying instruction tuning with Romanian datasets, such as RoAlpaca Taori et al. (2023); Masala et al. (2024), RoOpenAssistant Köpf et al. (2023); Masala et al. (2024) and RoOrca Mukherjee et al. (2023); Masala et al. (2024). The two LLMs have been optimized for Romanian morphology and syntax using tokenizers adapted to the language. They have shown strong performance on Romanian natural language understanding and generation tasks Masala et al. (2024). In our study, we further fine-tune both models on RoMedQA and compare their performance under different prompt structures and lengths.
RoLLaMA2-7B-Instruct. As it inherits the architectural constraints of the original LLaMA 2 model Touvron et al. (2023), RoLLaMA2-7B is limited to a context window of 4,096 tokens. We employ a LoRA configuration specialized on the causal language modeling task. The chosen configuration enables updating a small subset of the model parameters ($0.0622\%$) during fine-tuning, significantly reducing training overhead. To enhance the adaptability of the LLM to the clinical QA task, we further consider several configuration alternatives. First, we evaluate on various prompt lengths, considering limits of 1,024, 2,048 and 4,096 tokens. If an epicrisis is longer than the accepted input length, we simply trim it from the end. When evaluating on prompts of 4,096 tokens, we also explore two prompt formats: Q+E+A (question + epicrisis + answer) and E+Q+A (epicrisis + question + answer). The model is trained via token-level cross-entropy, where the task is to predict the answer tokens via next token prediction. For this reason, the answer is always placed at the end of the prompt.
RoMistral-7B-Instruct. Given the large number of parameters in the base Mistral model, we again employ a LoRA-based fine-tuning strategy, enabling updates to only a small number of parameters. In this case, we fine-tune just $0.0470\%$ of the parameters, using the same LoRA configuration as for RoLLaMA2-7B-Instruct. This approach ensures both memory efficiency and effective domain adaptation to the Romanian clinical QA setting. We conduct experiments with two prompt lengths, namely 2,048 and 4,096. Consistent with the preliminary empirical results obtained with the previous model, the prompt format is set to Q+E+A in all the experiments performed with RoMistral-7B-Instruct. We employ the token-level cross-entropy to optimize the model.
4.2 Long-Context LLM
Phi-4-mini-instruct is a compact instruction-tuned transformer model designed for long-context understanding, offering efficient inference for up to 128K tokens. We employ this model primarily due to its ability to process significantly longer input sequences than RoLLaMA2-7B-Instruct and RoMistral-7B-Instruct, respectively. Given that a substantial proportion of the epicrises exceed the 4,096-token limit (see Table 3), we hypothesize that Phi-4-mini-instruct could offer competitive performance by leveraging a broader context, despite being a smaller model, and without explicit pretraining on the Romanian language.
To adapt the model to the clinical QA setting, we fine-tune it using LoRA, updating only $0.0956\%$ of the parameters via token-level cross-entropy. We employ the same LoRA configuration as for RoLLaMA2-7B-Instruct and RoMistral-7B-Instruct. Thanks to the architectural support of Phi-4 for extended input lengths, we evaluate the model under increasingly longer prompts, namely 2,048, 3,072, 4,096, 8,192, and 16,384 tokens. This allows us to assess the relationship between available context and model performance, particularly in scenarios where clinical answers are distributed across epicrises of patients with a long medical history. All prompts follow the Q+E+A format.
4.3 Biomedical LLM
LLaMA3-OpenBioLLM-8B is a domain-specialized model designed to address biomedical and clinical tasks. It is based on the LLaMA3 architecture and has undergone extensive pretraining on a diverse collection of biomedical corpora, including high-quality resources such as PubMed abstracts, full-text clinical trials and biomedical textbooks. Through this targeted pretraining, the model achieves a strong understanding of medical terminology, domain-specific reasoning and structured clinical knowledge.
Despite being pretrained primarily in English, LLaMA3-OpenBioLLM-8B can still offer a valuable perspective when evaluated on Romanian clinical question answering. To adapt the model to our Romanian QA task, we train the model via LoRA (with the same configuration as for the other LLMs), updating only 3,407,872 parameters, which corresponds to $0.0424\%$ of the full model size. We use the Q+E+A prompt format and explore two different prompt lengths, 2,048 and 4,096.
4.4 Baselines
To assess the effectiveness of our fine-tuned models and to better isolate the impact of domain and task adaptation, we include three baselines.
Random token selector. For this baseline, we run three versions with different input lengths (1,024, 2,048, and 4,096 tokens). For each setup, we randomly select the starting token index (between 0 and the maximum input length) and a span length (between 1 and 3 tokens) to serve as the predicted answer. The answer is extracted from the input context (epicrisis), while entirely ignoring the question. This baseline serves as a trivial reference point, highlighting the complexity of the clinical QA task and providing a lower-bound performance estimate.
Majority answer model. Another simple baseline is based on answering with the most prevalent answer in the training set. This model consistently answers with “Nu” (“No”), regardless of the posed question or the epicrisis content. While still trivial, the majority answer model can be a more competitive baseline than the random token selector.
Zero-shot inference. For each of the evaluated models (RoLLaMA2-7B-Instruct, RoMistral-7B-Instruct, Phi-4-mini-instruct, and LLaMA3-OpenBioLLM-8B), we assess performance without any task-specific fine-tuning. We use the Q+E+A prompt format across all models. The zero-shot setup is meant to reflect the out-of-the-box capabilities of distinct types of LLMs.
Model Prompt format Validation Test F1 EM BLEU METEOR F1 EM BLEU METEOR RoLLaMA2-7B-Instruct E+Q+A 0.2518 0.2395 0.2233 0.1656 0.2871 0.2800 0.2531 0.1867 Q+E+A 0.2763 0.2545 0.2508 0.1452 0.3142 0.2890 0.2772 0.1647
Table 4: Preliminary results with alternative prompt formats for RoLLaMA based on 4,096 tokens. We compare two prompt formats: epicrisis + question + answer (E+Q+A) vs. question + epicrisis + answer (Q+E+A). The Q+E+A format yields higher performance across most evaluation metrics. The best results are highlighted in bold green.
5 Experiments
We conduct experiments with the chosen methods on the RoMedQA dataset, using our official split. We start by tuning various hyperparameters involved in the fine-tuning process, aiming to find the optimal hyperparameter setup for all LLMs. We next evaluate two different prompt formats, to determine which prompt format leads to higher performance. For the best performing prompt structure, we conduct several experiments across multiple LLMs, using various prompt lengths. The objectives of our experiments are: (i) to capture how Romanian-adapted models compare to general-purpose and domain-specific alternatives, (ii) to determine how fine-tuning performs against zero-shot inference, (iii) to establish how fine-tuned or zero-shot LLMs compare against trivial baselines, and (iv) to investigate how the context length influences model performance, particularly for models with extended context capabilities.
5.1 Evaluation
We employ four alternative evaluation metrics.
F1 score. The first metric that we take into consideration is the F1 score, which measures precision and recall at the token level. It provides a balanced view of how well models identify certain parts of the answer.
Exact match. We also consider the exact match (EM) score, which represents the percentage of predictions that perfectly match the ground-truth answer. Although this metric is the most demanding, it is crucial in medical contexts where small wording differences can change meaning.
BLEU. Because RoMedQA contains answers with more than one token, we consider the BLEU score Papineni et al. (2002), which measures n-gram overlap to assess fluency and lexical similarity. When the predicted answer includes only a subset of the tokens from the ground-truth answer, BLEU can provide a representative measure by considering partial n-gram matches.
METEOR. Lastly, we include METEOR Banerjee and Lavie (2005), which offers a more flexible evaluation by accounting for synonyms, stemming, and word order variations. These aspects are particularly valuable given the rich morphology and flexible syntax of Romanian.
Model Fine-tuned #Tokens Validation Test F1 EM BLEU METEOR F1 EM BLEU METEOR Random token selector N/A 1,024 0.0023 0.0006 0.0006 0.0013 0.0019 0.0001 0.0004 0.0011 2,048 0.0021 0.0004 0.0004 0.0013 0.0024 0.0003 0.0005 0.0015 4,096 0.0022 0.0005 0.0006 0.0014 0.0026 0.0003 0.0005 0.0015 Majority answer N/A N/A 0.2091 0.1786 0.1831 0.1153 0.2148 0.1838 0.1884 0.1183 RoLLaMA2-7B-Instruct 2,048 0.0188 0.0015 0.0024 0.0146 0.0230 0.0025 0.0023 0.0166 1,024 0.5535 0.5431 0.4996 0.3498 0.5646 0.5300 0.4896 0.3451 2,048 0.5832 0.6035 0.5651 0.3487 0.5944 0.6145 0.5780 0.3554 4,096 0.2763 0.2545 0.2508 0.1452 0.3142 0.2890 0.2772 0.1647 RoMistral-7B-Instruct 4,096 0.0715 0.0430 0.0234 0.0438 0.0605 0.0350 0.0229 0.0403 2,048 0.5761 0.5820 0.5387 0.3425 0.5808 0.5800 0.5331 0.3477 4,096 0.5814 0.5925 0.5487 0.3456 0.5963 0.6045 0.5546 0.3590 Phi-4-mini-instruct 2,048 0.0078 0.0000 0.0007 0.0075 0.0048 0.0002 0.0005 0.0043 2,048 0.6500 0.6846 0.6366 0.3916 0.6618 0.6935 0.6468 0.3976 3,072 0.6617 0.6960 0.6472 0.3974 0.6671 0.7011 0.6522 0.4009 4,096 0.4217 0.4145 0.3923 0.2274 0.4220 0.4155 0.3799 0.2303 8,192 0.5050 0.4960 0.4624 0.2911 0.5216 0.5105 0.4653 0.3037 16,384 0.5168 0.5110 0.4765 0.2991 0.5267 0.5130 0.4731 0.3060 LLaMA3-OpenBioLLM-8B 2,048 0.0140 0.0004 0.0014 0.0156 0.0119 0.0004 0.0015 0.0132 2,048 0.5861 0.6140 0.5720 0.3498 0.5958 0.6215 0.5792 0.3585 4,096 0.5365 0.5855 0.5210 0.3163 0.5566 0.5775 0.5382 0.3348
Table 5: Results of multiple LLMs based on various configurations vs. two trivial baselines on RoMedQA. For each LLM, there is a zero-shot version and several fine-tuned versions, with various prompt lengths. F1, Exact Match (EM), BLEU, and METEOR scores are reported on both validation and test sets. Fine-tuned models consistently outperform zero-shot versions. The best results for each LLM are highlighted in bold green.
5.2 Preliminary Setup and Experiments
Hyperparameter tuning. We employ consistent hyperparameter settings across all models to ensure a fair comparison among LLM families. All models are fine-tuned using the AdamW Loshchilov and Hutter (2019) optimizer. Due to memory constraints, we set a per-device batch size of one sample with gradient accumulation over $8$ steps, resulting in an effective mini-batch size of $8$ samples. Training is carried out for $5$ epochs, using Brain Floating Point (BFLOAT16) precision to optimize memory usage, while maintaining numerical stability. The learning rate, the dropout rate and the LoRA configuration are established via grid search on the validation set. The range for the initial learning rate is $10^{-3}$ to $10^{-6}$. We employ a cosine learning rate scheduler with $100$ warm-up steps. Dropout is applied with drop rates between 0 and $0.2$, using a step of $0.05$. For the LoRA rank $r$, we look for an optimal value in the set $\{4,6,8,10\}$. For the LoRA scaling factor, we consider values in the set $\{8,16,32\}$. The optimal configuration is based on a learning rate of $2\cdot 10^{-5}$, a dropout rate of $0.05$, a LoRA rank $r$ of $8$, and a LoRA scaling factor $\alpha$ of $16$.
Model Text chunks Validation Test F1 EM BLEU METEOR F1 EM BLEU METEOR Phi-4-mini-instruct First 0.6500 0.6846 0.6366 0.3916 0.6618 0.6935 0.6468 0.3976 Phi-4-mini-instruct All 0.4799 0.5170 0.4106 0.2934 0.4913 0.5288 0.4191 0.2989
Table 6: Results of Phi-4-mini-instruct on RoMedQA, using only the first 2,048-token chunk of the epicrisis vs. all 2,048-token chunks. The former approach significantly outperforms the latter one across all metrics, suggesting that only analyzing the first part of the epicrisis reduces input clutter and prevents overfitting. The best results are highlighted in bold green.
Figure 3: Performance per question category (binary, extractive, and reasoning) across four evaluation metrics (F1, EM, BLEU, and METEOR). The values are reported for the top scoring model, namely Phi-4-mini-instruct based on 3,072 tokens, on the official test set. Best viewed in color.
Question (Romanian / English) Ground-truth Predicted Correct?
Imunoterapia a fost introdusă în a doua etapă a schemei terapeutice? / Was immunotherapy introduced in the second stage of the therapeutic regimen? Nu / No Nu / No
S-au identificat metastaze la nivelul plămânilor? / Are lung metastases present? Da / Yes Da / Yes
Care este stadiul tumoral la diagnostic al pacientului? / What was the tumor stage at the time of diagnosis for this patient? IIA IIA
Care este subtipul histologic al pacientului din rezultatul histopatologic? / What is the histological subtype of the patient according to the histopathology report? ADK ADK
Ce cadrane ale sânului sunt invadate de tumoră? / Which breast quadrants are invaded by the tumor? supero-extern / upper-outer supero-extern / upper-outer
S-au utilizat alte linii de tratament în progresia bolii? / Were other lines of treatment used during disease progression? CHT Nu / No
S-au identificat metastaze la nivelul oaselor? / Are bone metastases present? Nu / No Da / Yes
Sindrom inflamator este prezent la acest pacient? / Is the inflammatory syndrome present for this patient? prezent / present absent / absent
Care este stadializarea TNM a pacientului? / What is the TNM staging of the patient? cT4cT2cN2M0 / pT4bPNx cT4cN2M0
Care este stadiul tumorii stabilit la începutul tratamentului? / What is the stage of the tumor established at the beginning of treatment? IV IIIB
Table 7: Examples of test questions with correct and wrong answers provided by the best performing LLM, namely Phi-4-mini-instruct based on 3,072 tokens. Each row shows a question along with ground-truth and predicted answers. English translations are provided when necessary.
Prompt format. We conduct a preliminary experiment to assess the impact of the prompt structure on QA performance. Specifically, we compare two input formulations for the RoLLaMA2-7B-Instruct model: epicrisis + question + answer (E+Q+A) vs. question + epicrisis + answer (Q+E+A). Both configurations are evaluated with 4,096 input tokens. As shown in Table 4, the Q+E+A format consistently outperforms the E+Q+A format across all evaluation metrics, besides METEOR. Given these results, we decided to run all subsequent experiments using the Q+E+A prompt structure.
5.3 Results
In Table 5, we present the results obtained by the chosen LLMs vs. two baselines on RoMedQA. We evaluate all four LLMs under both zero-shot prompting and supervised fine-tuning setups, reporting performance metrics on both validation and test sets.
Baselines vs. LLMs. In the zero-shot setup, the performance levels reached by the various LLMs indicate that out-of-the-box models can surpass the naive random token selector. However, the majority answer baseline, which leverages knowledge about the answer distribution in training data, significantly outperforms all zero-shot variants. This is a first hint that zero-shot models are not able to generalize to the medical QA task in Romanian, regardless of their prior language or domain adaptation. In contrast, each fine-tuned LLM version surpasses the two baseline models, irrespective of prompt length and prior model adaption. This observation holds for all LLM architectures, indicating that the fine-tuning process is generally useful.
Zero-shot vs. fine-tuned LLMs. In general, zero-shot models exhibit significantly lower performance than their fine-tuned counterparts. Furthermore, for various performance metrics and LLM architectures, we observe that the scores can increase by an order of magnitude when going from zero-shot prompting to fine-tuning. The comparison between zero-shot and fine-tuning setups highlights the critical importance of task-specific fine-tuning for clinical QA, confirming that prior language-specific or domain-specific adaptation is insufficient to accurately address the task.
Effect of prompt length. We test the models with various prompt lengths, taking into account the limitation imposed by each LLM architecture on its input prompt. While the average length of an epicrisis is 7,171 tokens (as per Table 3), the LLMs do not necessarily benefit from using extensive prompts. For example, RoLLaMA2-7B-Instruct, Phi-4-mini-instruct and LLaMA3-OpenBioLLM-8B obtain superior performance levels with 2,048 input tokens than with 4,096 tokens. The only model that obtains better results with 4,096 tokens instead of 2,048 is RoMistral-7B-Instruct, but the performance gains are rather marginal.
Phi-4-mini-instruct is the only model that accepts very long prompts. Although we explore prompts of up to 16,384 tokens for Phi-4-mini-instruct, the model does not benefit from the extended input.
First vs. multiple input chunks. Since epicrises are trimmed from the end to fit the desired prompt lengths, the generally low performance levels when using long prompts can be explained by two factors: (i) most questions can be answered just by looking at the beginning part of an epicrisis, or (ii) taking long contexts into account is a challenging learning task, potentially leading to overfitting. To distinguish between these two cases, we conduct another experiment with one of the best performing models, Phi-4-mini-instruct based on 2,048 input tokens. In addition to the version that simply trims the epicrisis from the end, we evaluate a version that analyzes non-overlapping 2,048-token chunks from the epicrisis. For the latter approach, we employ majority voting over all chunks to obtain the final answer. To break ties, we rely on the average log-probability over the generated tokens and select the answer with the highest value. The corresponding results are reported in Table 6. While the chunk-based approach enables processing of complete epicrises, it results in substantial performance degradation, suggesting that focusing on the first part of the epicrisis leads to optimal results. Still, as shown by our previous experiments reported in Table 5, using a context that is too short, e.g. 1,024 tokens for RoLLaMA2-7B-Instruct or 2,048 tokens for RoMistral-7B-Instruct, might also lead to performance degradation. In summary, trimming the end part of the epicrisis seems to act as a regularization technique, and finding the optimal context length can be assimilated with the process of tuning the corresponding regularization hyperparameter.
Use of multiple evaluation metrics. Regarding the employed evaluation metrics, we observe that EM scores exceed F1 scores by small margins across most configurations, suggesting that the models tend to either match answers exactly or miss them entirely, with fewer partial matches. Moreover, METEOR scores are consistently lower than other metrics, likely reflecting the sensitivity of METEOR to synonyms and word variations in Romanian medical terminology, but also its inherent bias against short answers. Nevertheless, for a holistic evaluation of QA models on RoMedQA, we promote the use of all four evaluation metrics.
Performance per question type. In Figure 3, we present the results of our best performing model, for each of the three question categories. The F1 and EM metrics suggest that binary questions are easier to answer, while reasoning questions are harder. This observation is consistent with our natural expectation. However, the BLEU score seems to promote reasoning questions, while METEOR indicates that the model is more capable on extractive questions. Overall, the analysis shows that some metrics are more appropriate for certain question categories. This represents another strong argument in favor of using a wide variety of metrics to accurately benchmark models on RoMedQA.
General remarks. Our experimental results reveal several patterns across all models and configurations. First, we observe that validation and test results remain consistent across all experiments, indicating that our patient-level data splits effectively prevent data leakage. All fine-tuned LLMs exceed the naive baselines by large margins, confirming that models learn meaningful clinical question-answering capabilities. We also find that fine-tuned models significantly outperform their zero-shot counterparts across all architectures, emphasizing that neither prior Romanian language adaptation nor English biomedical pretraining are sufficient for the medical QA task. Among the four LLM families, RoMistral-7B-Instruct obtains the lowest performance levels. In contrast, Phi-4-mini-instruct achieves the highest scores among all evaluated models, despite not being explicitly pretrained on the Romanian language nor on the medical domain. These results confirm our previous observation and similar studies Dorfner et al. (2024), suggesting that prior language or domain adaptation is insufficient.
5.4 Qualitative Analysis
To better understand the challenges posed by RoMedQA and the capabilities of the best-performing model, we conduct a qualitative analysis of a random selection of QA pairs with correct and wrong answers predicted by Phi-4-mini-instruct. In Table 7, we showcase 10 examples, 5 that are correctly answered and 5 that are incorrectly answered, respectively.
Analysis of correctly answered questions. In the first example, the LLM was able to recognize the sequence of treatment lines and to distinguish between the first and the second line of therapy, even though the information was not explicitly stated in the epicrisis. In the second example, pulmonary metastases were mentioned under the abbreviation M1PUL. The model recognized that M1PUL denotes the presence of lung metastases, providing the correct answer based on this information. In the third example, the LLM demonstrates the ability to correctly stage breast cancer, which requires high-level medical reasoning. The model correlated tumor size, regional invasion, and the absence of distant metastases, successfully assigning stage IIA, according to the TNM criteria. In the fourth case, the histological type “adenocarcinom” / “adenocarcinoma” was explicitly stated in the histopathology report, which allowed the LLM to correctly answer with “ADK” to the posed question. In the fifth example, the interpretation of the chest CT explicitly stated “invazia tumorală în cadranul supero-extern” / “tumor invasion present in the upper-outer quadrant”. The model was able to accurately spot this information and provide the correct answer. Overall, we observe that the model demonstrates the necessary capabilities to answer both reasoning and extractive questions.
Analysis of wrongly answered questions. In the first example, the annotator considered that the presence of “CHT” in the epicrisis indicates a new line of treatment at disease progression. However, the LLM answered “Nu” / “No”, either because it did not recognize that “CHT” represented a new line of therapy, or because it required an explicit statement confirming the change of treatment line, which is missing from the epicrisis. In the latter case, the model might have considered that “CHT” is part of the initial treatment, instead of a new line of therapy. In the second example, the LLM may have been triggered by the presence of the word “os” / “bone”. The confusion is caused by an investigation reported in the epicrisis, called “scintigrafie osoasă” / “bone scintigraphy”, which is usually recommended for patients with bone metastases. Despite undergoing this investigation, the patient did not have any bone metastases. In the third case, the inflammatory syndrome is not directly mentioned in the epicrisis, although it can be identified by analyzing biological parameters (e.g. elevated ESR, increased CRP, elevated fibrinogen) or secondary diagnoses (e.g. infection, chronic inflammation). The physician recognized these signs and determined the presence of the inflammatory syndrome. Without an explicit mention of “sindrom inflamator” / “inflammatory syndrome”, the LLM was not able to extract the correct answer. In the fourth case, the LLM provided only a partially correct answer. It correctly identified the clinical TNM classification, but did not mention the pathological TNM classification. This is a limitation induced by the prompt length of 3,072 tokens, since the pathological TNM classification is usually specified near the end of the epicrisis, which gets removed due to context trimming. In the fifth example, the patient had lymph node metastases. The physician considered the staging to be stage IV, whereas the LLM interpreted the involvement as loco-regional lymph node invasion and classified the case as stage IIIB. Overall, the wrong or partially incorrect answers point out two kinds of limitations: (i) the reasoning abilities of Phi-4-mini-instruct can sometimes prevent it from integrating high-level medical information, and (ii) the context trimming procedure might sometimes lead to dropping important clues.
6 Conclusion
In this paper, we introduced RoMedQA, the first benchmark for Romanian medical question answering. The new benchmark is the result of a meticulous data collection and annotation process carried out by experts in the medical domain, who constructed a dataset of 102,646 QA pairs. The QA samples are grounded in a collection of 1,011 medical case summaries of cancer patients. We conducted comprehensive experiments with several LLM families to assess the impact of various aspects, namely prompt structure and length, language vs. domain vs. general pretraining, and zero-shot prompting vs. LoRA-based fine-tuning. Our empirical results suggested some interesting findings. First of all, we observed that long prompts are not always helpful, degrading performance even if epicrises are typically long documents. Second of all, we found that generic pretraining can lead to better results, suggesting that prior language or domain adaption can actually produce LLMs with reduced generalization capabilities. Third of all, we determined that fine-tuning on our specific task significantly outperforms zero-shot prompting. Yet, even the best model was only able to correctly answer about $70\%$ of the test questions, indicating that RoMedQA is a challenging benchmark.
In future work, we aim to develop models that integrate retrieval-augmented generation to improve question answering capabilities. More specifically, we will focus on developing a retrieval module to accurately find the sections of the input epicrisis that are likely to contain indicators for the answer. This will allow us to use models with reasonable context lengths, while minimizing the risk of eliminating important information during context trimming. However, since the answer is not always directly mentioned in the epicrisis, finding the right sections from which the answer could be derived requires careful consideration.
Acknowledgments
This research is supported by the project “Romanian Hub for Artificial Intelligence - HRIA”, Smart Growth, Digitization and Financial Instruments Program, 2021-2027, MySMIS no. 334906.